OpenStudy (anonymous):

Verify: (sec(m)+1)/(tan(m))=csc(m)+cot(m)

zepdrix (zepdrix):

\[\Large\rm \frac{\sec m+1}{\tan m}\]You can split up the fractions like this, \[\Large\rm =\frac{\sec m}{\tan m}+\frac{1}{\tan m}\]Understand that step?

OpenStudy (anonymous):

i figured out that question. I need help on a new one now verify the following: ((sin2x)/cot(x))=1-cot(2x)